OpenStudy (mathhelpplease):

I need physics help. If you could walk me through this, Id be very appreciative of it. If a ball is thrown off of a 123 m cliff at a speed of 18 m/s, how far from the base of the cliff is the ball?

OpenStudy (mathmale):

In which direction is the ball thrown? Straight down? Horizontally? At an angle? Note that if the ball is thrown straight down from the top of the cliff, it will theoretically land at the base of the cliff. That sounds too easy to be correct. If the ball is thrown away from the top of the cliff at a 45 degree angle from the horizontal, then the ball will hit the ground some distance from the base of the cliff, and you can determine how far. Are you sure you've shared every detail from the original problem statement?

OpenStudy (kevin):

First, we find how much time (t) that the ball need for throw off and touch the ground : -y = voy t - 1/2gt^2 -123 = 0 - 1/2 x 10 x t^2 t = 4,96 s Now we find the horizontal distance (x) between base of the cliff and the ball x = Vox x t x = 18 x 4.96 x = 89.28 m